OpenStudy (anonymous):

Solve the compound inequality. 5x + 11 ≥ –9 and 10x – 3 ≤ 27 A. x ≥ –4 and x ≤ 3 B. x ≥ –4 and x ≤ 2 2/5 C. x ≥ 2/5 or x ≤ 2 2/5 D. x ≥ 2/5 or x ≤ 3

OpenStudy (nubeer):

for 5x+11 >= -9 subtract 11 on both sides.. then divide both sides by 11 what will you get?

OpenStudy (anonymous):

5x > -20

OpenStudy (anonymous):

He meant divide both sides by 5 to find x.

OpenStudy (anonymous):

ah that makes more sense

OpenStudy (anonymous):

x > -4

OpenStudy (anonymous):

Yes. Then solve for x in the other expression: 10x-3 <=27

OpenStudy (anonymous):

Add 3 to both sides. The divide both sides by 10 to find x.

OpenStudy (anonymous):

x ≤ 3

OpenStudy (anonymous):

Yes ans.: x>=-4 and x<=3. The 'and' comes from the 'and' in the original question.

OpenStudy (anonymous):

So that would be A.

OpenStudy (anonymous):

Got any questions?

OpenStudy (anonymous):

thanks. Do you think you could help me on another question?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Solve the absolute value equation. │3x + 1│ = 1

OpenStudy (anonymous):

1. To solve an absolute value equality question the absolute expression needs to be by itself on one side: |3x+1| = 1. So we have that. 2. The absolute value of an expression is always greater than or equal to zero. If the number on the right is negative then there is "NO SOLUTION". In this case there is a solution because we have a positive '1'. 3. We have 2 cases: 3x+1 = 1 or 3x+1=-1 Then solve for x in both cases.

OpenStudy (anonymous):

i know from the options its either B or D. B. x = -2/3 or x = 0 D. x = 1 or x = 0

OpenStudy (anonymous):

3x+1 = 1 or 3x+1=-1 Solve for x in both equations above, what do you get?

OpenStudy (anonymous):

B. x = -2/3 or x = 0

OpenStudy (anonymous):

Good deal, any questions about how i got this?

OpenStudy (anonymous):

Nope. You've explained pretty well

OpenStudy (anonymous):

Do you have anymore questions you would like to ask?

OpenStudy (anonymous):

probably but not at the moment

OpenStudy (anonymous):

@philo1234 How do you figure out sloe intercepts?

OpenStudy (anonymous):

x-intercept is where the graph crosses the x-axis at this point y =0 (always). y-intercept is where the graph crosses the y-axis, at this point x=0 (always). Eg.: y = x + 2. The x-int is x+2=0. That is x=-2. The y-int is y=0+2. That is y=2

OpenStudy (anonymous):

so for –6x + 3y = 12. it would be: -6x + 3(0) = 12. -6x = 12. x = -2. -6(0) + 3y = 12. 3y = 12. y = 4.

OpenStudy (anonymous):

Yes that's correct

OpenStudy (anonymous):

Yay =^..^=

OpenStudy (anonymous):

how do I figure out the equation based on the slope?

OpenStudy (anonymous):

The questions come in 2 forms: 1. Give you 2 points on a line: a. Find the slope using the 2 points. b. Use the point slope form: y-y1=m(x-x1), where x1 and y1 is the x and y value of one of the points and m is the slope. 2. Give you the slope and a point: a. Use the point slope form: y-y1=m(x-x1), where x1 and y1 is the x and y value of the point given and m is the slope..

OpenStudy (anonymous):

the problem only gives me point: (9, 3); slope: 4/9

OpenStudy (anonymous):

Eg1. Find the equation of the line given the point (3,0) and (1,4). Slope (m) = -2 Choose one of the points: (3,0) - x1 = 3 and y1=0 Point-slope form: y-0=-2(x-3) Answer in slope intercept form: y=-2x+6 Answer in standard form (ax+by=c) = 2x+y=6 Eg2. Find the equation of the line given the slope -2 and the point (1,4) m=-2, x1=1, y1=4. Point-slope form: y-4=-2(x-1) Answer in slope intercept form: y=-2x+6 Answer in standard form (ax+by=c) = 2x+y=6

OpenStudy (anonymous):

So your question is similar to Eg2 use the same approach to find the equation. Can you do that?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

y - 9 = 4/9(x-3)

OpenStudy (anonymous):

nope: y-3=4/9(x-9), remember point-slope form: y-y1=m(x-x1). Should your answer be in a specific form?

OpenStudy (anonymous):

that is the form it wants

OpenStudy (anonymous):

So answer would be: y-3=4/9(x-9), Note that the 3 is the y-value and 9 is the x-value taken from the point given.